Assuming all sex distributions to be equally probable, what proportion of families with exactly six children should be expected to have three boys and three girls?
Answer by stanbon(75887)

Assuming all sex distributions to be equally probable, what proportion of families with exactly six children should be expected to have three boys and three girls?
----
# of 6-child families with 3 boys = 6C3 = 20
---
# of different boy-girl family patterns: 2^6 = 64
-------
Proportion: 20/64 = 5/16
============================
Cheers,
Stan H.